How many 5 letter code words can be formed from the letters in OUTFIELD if no letter is repeated and the last letter is a vowel? I don't know how to solve the problem and making the last letter a vowel, it's been irking me for a while. Any help will be appreciated!
Answer by stanbon(75887)

How many 5 letter code words can be formed from the letters in OUTFIELD if no letter is repeated and the last letter is a vowel?
8 letters
4 vowels
--------------
Choose the last letter---4 ways
Choose the other 4 letters without repetition ---7*6*5*4 ways
=============
Ans: 4(7*6*5*4)= 4*840= 3360 code words
============
Cheers,
Stan H.